5/4x + 1/8= 7/8 solve the solution is x=
I came up with subtract 1/8 from both sides = 7/8 -1/8= 6/8 reduced x= 1/2 am I right

Found 2 solutions by solver91311, nyc_function:
Answer by solver91311(24713)


You were on the right track but got derailed somewhere in the middle.



Adding to both sides was the correct thing to do, but you didn't finish the simplification.







Now I don't know where you went. You have a coefficient on , so you need to multiply both sides by the reciprocal of that coefficient:



And simplify:




John


Answer by nyc_function(2741)

Did you check your answer?
You said x = 1/2
Plug x = 1/2 into your orginal question and simplify. If you get the same answer on both sides, then you will know that the answer you found is right.
In other words, if you don't get 7/8 = 7/8, then your answer is not right.
Do the checking and get back to me.